OpenStudy (anonymous):

The given equation is either linear or equivalent to a linear equation. Solve the equation. (If there is no solution, enter NO SOLUTION. If all real numbers are solutions, enter REALS.) Z/5 = 3/10 z + 8

OpenStudy (dumbcow):

cross multiply z*(10z+8) = 15 distribute 10z^2 + 8z = 15 set equal to 0 10z^2 + 8z -15 = 0 use quadratic formula -8 +-sqrt(64-4(10)(-15)) / 20 = (-8 +- 25.77)/20 =.88 or -1.68

OpenStudy (anonymous):

shouldnt it = a fraction

OpenStudy (dumbcow):

i just realized i did this wrong, i noticed its supposed to be a linear function huh? so the equation is probably 1/5z = 3/10z + 8 get z's on one side -1/10z = 8 multiply by -10 z = -80